# Written by AAAI Press Staff1

### v1

@@ 5 Conclusion
In this paper, we have introduced a novel graph reinforcement learning framework NeuroCUT to effectively address the graph partitioning problem. A distinctive feature of our framework is its adaptability to diverse partitioning objectives that may or may not be differentiable, which sets it apart from conventional algorithms. Our approach has also shown the ability to generalize on unseen graphs as well as unseen number of partitions. The efficacy of our approach is empirically validated through an extensive evaluation on six datasets with four graph partitioning objectives. Notably, our method shows remarkable performance when compared to both non-neural and neural state-of-the-art techniques, proving its competence in both inductive and transductive settings. In summary, NeuroCUT offers a versatile, inductive, and empirically validated solution for graph partitioning.
Dataset $|\mathcal{V}|$ $|\mathcal{E}|$ Method $normalised$ $sparsest$ $balanced$ $kmin$ AVG DIFF MAX DIFF
ACTDIFF ACTDIFF ACTDIFF ACTDIFF
Cora 2485 5069 hMetis 0.343   0.115 1.412   0.061 0.345   0.000 0.069   0.063 0.060 0.115
Spectral 0.370   0.142 1.351   0.000 1.165   0.820 0.006   0.000 0.240 0.820
GAP 0.505   0.277 2.074   0.723 0.536   0.191 0.095   0.089 0.320 0.723
NeuroCUT 0.228   0.000 1.351   0.000 0.570   0.225 0.011   0.005 0.057 0.225
CiteSeer 2120 3679 hMetis 0.198   0.025 0.673   0.227 0.198   0.000 0.038   0.037 0.072 0.227
Spectral 0.173   0.000 0.446   0.000 0.929   0.731 0.001   0.000 0.182 0.731
GAP 0.267   0.094 0.795   0.349 0.312   0.114 0.043   0.042 0.150 0.349
NeuroCUT 0.211   0.038 0.712   0.266 0.296   0.098 0.036   0.035 0.109 0.266
SBMs 500 5150 hMetis 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
Spectral 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
GAP - - - 0.035   $0$
NeuroCUT 0.191   0.000 3.939   0.000 0.191   0.000 0.038   0.003 0.0008 0.003
Facebook 1034 26749 hMetis 1.115   0.858 52.319   48.198 1.115   0.143 0.200   0.197 12.349 48.198
Spectral 1.670   1.413 4.722   0.601 2.459   1.487 0.003   0.000 0.875 1.487
GAP - - - 0.025   0.022 - -
NeuroCUT 0.257   0.000 4.121   0.000 0.972   0.000 0.015   0.012 0.003 0.012
Harbin 6235 15205 hMetis 0.058 0.000 0.184   0.000 0.058   0.000 0.011   0.000 0.000 0.000
Spectral 0.104   0.046 0.348   0.164 0.403   0.345 0.015   0.004 0.140 0.345
GAP 0.175   0.117 0.551   0.367 0.269   0.211 0.017   0.006 0.175 0.367
NeuroCUT 0.092   0.034 0.286   0.102 0.099   0.041 0.018   0.007 0.046 0.102
Table 7: Results of all methods on all five datasets. “ACT” denotes the actual values of the measures and “DIFF” shows the difference with the best value for the dataset. “AVG DIFF” and “MAX DIFF” present the average and the maximum of the differences over all the measures respectively. The best average differences are shown by green and the second best ones are in gray. The worst result, i.e., the maximum difference is shown in red. The neural state-of-the-art GAP method produces the most inferior performance among considered baselines. Our method NeuroCUT either outperforms the baselines (green) or produces the second best (gray) results while being close to the best one in all cases. Moreover, NeuroCUT is the most robust and never produces the worst maximum difference (i.e., in red). The neural methods are trained and tested with the same number partitions ($k=5$).
Appendix A Result for all datasets in transductive setting
In Table  7 we present the results obtained by all methods on different datasets. These correspond to radar plot obtained in fig. 2 in main paper .
Appendix B Clustering for Initialization
@@
As discussed in sec 3.1 in main paper, we first cluster the nodes of the graph into $k$ clusters where $k$ is the number of partitions. Towards this, we apply K-means algorithm (macqueen1967some) on the nodes of the graph where a node is represented by its positional node embedding i.e $pos(u)\;\forall u\in\mathcal{V}$. Further, we used $L_{\inf}$ norm as the distance metric for clustering.
Appendix C Time complexity of NeuroCUT
@@
(1) First the positional embeddings for all nodes in the graph are computed. This involves running RWR for $\alpha$ anchor nodes for $\beta$ iterations(Eq.  6). This takes $\mathcal{O}(\alpha\times\beta)$ time. (2)Next, GNN is called to compute embeddings of node. In each layer of GNN a node $v\in\mathcal{V}$ aggregates message from $d$ neighbors where $d$ is the average degree of a node. This takes $\mathcal{O}(\mathcal{V})$ time. This operation is repeat for $L$ layers. Since $L$ is typically 1 or 2 hence we ignore this factor. (3) The node selection algorithm is used that computes score for each node based upon its neighborhood using eq. 9. This consumes $\mathcal{O}(|\mathcal{V}|\times d)$ time. (4) Finally for the selected node, its partition has to be determined using eq. 11 and  12. This takes $\mathcal{O}(d+k)$ time, as we consider only the neighbors of the selected node to compute partition score. Steps 2-4 are repeated for $T^{\prime}$ iterations. Hence overall running time is $\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{V}|\times d+k)\times T^{\prime}\big{)}$. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Since $|\mathcal{V}|\times d\approx|\mathcal{E}|$, hence complexity is $o(|\mathcal{E}|\times|\mathcal{V}|\big{)}$ Further, for sparse graphs $|\mathcal{E}|=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.
Appendix D Impact of number of anchor nodes
In this section we analyze the impact of number of anchor nodes chosen for computing positional embeddings of nodes in the graph $pos(u)$. Fig. 5 shows the results where $Y$-axis shows the normalized cut value and $X$-axis shows the number of anchor nodes. The best results are achieved at the number of anchor nodes being 35 except in Citeseer. The other variation in the number of anchor nodes produces similar results often.
Figure 5: Results on varying the number of anchor nodes on the normalized cut value produced by NeuroCUT.

### v2

@@ 5 Conclusion
In this work, we study the problem of graph partitioning with node features. Existing neural methods for addressing this problem require the target objective to be differentiable and necessitate prior knowledge of the number of partitions. In this paper, we introduced NeuroCUT, a framework to effectively address the graph partitioning problem with node features. NeuroCUT tackles these challenges using a reinforcement learning-based approach that can adapt to any target objective function. Further, attributed to its decoupled parameter space and partition count, NeuroCUT can generalize to an unseen number of partitions. The efficacy of our approach is empirically validated through an extensive evaluation on four datasets, four graph partitioning objectives and diverse partition counts. Notably, our method shows significant performance gains when compared to the state-of-the-art techniques, proving its competence in both inductive and transductive settings. In summary, NeuroCUT offers a versatile, inductive, and empirically validated solution for graph partitioning.
Appendix A Appendix
A.1 Additional Related Work
A.2 Clustering for Initialization
@@
As discussed in sec 3.1 in main paper, we first cluster the nodes of the graph into $k$ clusters where $k$ is the number of partitions. Towards this, we apply K-means algorithm (macqueen1967some) on the nodes of the graph where a node is represented by its raw node features and positional representation i.e $\mathbf{emb_{init}}(u)\;\forall u\in\mathcal{V}$ based upon eq. 8. Further, we used $L_{\inf}$ norm as the distance metric for clustering.
A.3 Dataset Statistics
Table  7 describes the statistics of datasets used in our work. For all datasets we use their largest connected component.
Dataset $|\mathcal{V}|$ $|\mathcal{E}|$ Number of features
Cora 2485 5069 1433
CiteSeer 2120 3679 3703
Harbin 6598 10492 2
Actor 6198 14879 931
Table 7: Dataset Statistics
A.4 Time complexity of NeuroCUT
@@
(1) First the positional embeddings for all nodes in the graph are computed. This involves running RWR for $\alpha$ anchor nodes for $\beta$ iterations (Eq.  6). This takes $\mathcal{O}(\alpha\times\beta)$ time. (2) Next, GNN is called to compute embeddings of node. In each layer of GNN a node $v\in\mathcal{V}$ aggregates message from $d$ neighbors where $d$ is the average degree of a node. This takes $\mathcal{O}(\mathcal{V})$ time. This operation is repeat for $L$ layers. Since $L$ is typically 1 or 2 hence we ignore this factor. (3) The node selection algorithm is used that computes score for each node based upon its neighborhood using eq. 10. This consumes $\mathcal{O}(|\mathcal{V}|\times d)$ time. (4) Finally for the selected node, its partition has to be determined using eq. 12 and  13. This takes $\mathcal{O}(d+k)$ time, as we consider only the neighbors of the selected node to compute partition score. Steps 2-4 are repeated for $T^{\prime}$ iterations. Hence overall running time is $\mathcal{O}\big{(}(\alpha\times\beta)+(|\mathcal{V}|\times d+k)\times T^{\prime}\big{)}$. Typically $\alpha$, $\beta$ and $k$ are $<<|\mathcal{V}|$ and $T^{\prime}=o(|\mathcal{V}|)$. Since $|\mathcal{V}|\times d\approx|\mathcal{E}|$, hence complexity is $o(|\mathcal{E}|\times|\mathcal{V}|\big{)}$ Further, for sparse graphs $|\mathcal{E}|=\mathcal{O}(|\mathcal{V}|)$. Hence time complexity of NeuroCUT is $o(|\mathcal{V}|^{2})$.